if a piece of work could be completed in 24 days by 98 men . how many men would taketo complete it in 21 days ?
Answer by lenny460(1073)

If a piece of work could be completed in 24 days by 98 men. How many men would take to complete it in 21 days ?

24 days = 98 men
21 days = x men
Cross multiply:
24 * x = 98 * 21
24x = 2058
Divide each side by 24
24x/24 = 2058/24
x = 85.75
There is no 0.75 (three quarters) of a man.
Therefore:
We round up x to 86 men

Answer:
86 men




Lennox Obuong
Algebra Tutor
Nairobi, Kenya